OpenStudy (anonymous):

32 + (8 - 2) • 4 - 6/3..... 6 is over 3

OpenStudy (anonymous):

Follow the order of operations. Evaluate the expression in the parentheses first. 32+(8-2)*4-6/3 =32+6*4-6/3 Now, do mult/division in the order they appear. 32+6*4-6/3 =32+24-2 Finally, do addition/subtraction in the order they appear. 32+24-2 =56-2 =54

OpenStudy (anonymous):

or did you mean the entire expression is over 3?

OpenStudy (anonymous):

no just 6

OpenStudy (anonymous):

k